OpenStudy (anonymous):

solve each system using elimination: 1. 3x+3y=27 x-3y=-11 2. -x+5y=13 x-y=15 3. 2x+4y=22 2x-2y=-8 thanksies

OpenStudy (zepp):

3x+3y=27 x-3y=-11 In this one, do you see something you can get rid of?

OpenStudy (zepp):

Here: http://puu.sh/qCxy

OpenStudy (anonymous):

OHHH okay

OpenStudy (anonymous):

I'm confused, what do I do with the bottom equation? I'm really lost with this.

OpenStudy (zepp):

In the second one, -x+5y=13 x-y=15 You can see that there's a x and a -x, a y and a -y, do can use either, but they, I would suggest to transform the -y in the 2nd equation to 5y, how do you do that!? :O Hint: -y(x) = -5y Solve for x

OpenStudy (zepp):

Here's the solution for the second one: http://puu.sh/qCzY Now try to do the third one by yourself, keep in mind to find two similar terms but with opposite sign, good luck ;D (P.s.: Don't forget to find y's value as well :P )

OpenStudy (anonymous):

Thanks :D